m=10 kg
d=0.5 m
F=100N
f=100rd
t=?
1
Expert's answer
2019-12-13T10:08:01-0500
"a=\frac{F}{m}=\frac{100}{10}=10\frac{m}{s^2}"

"d=0.5at^2"

"0.5=0.5(10)t^2"

"t=0.316\ s"